OpenStudy (anonymous):

For My Little Sisters Thanks (; What is the area of a circle with diameter 10 inches 25 pi in2 20 pi in 2 5 pi in 2 100 pi in 2

jimthompson5910 (jim_thompson5910):

Area = pi * (radius)^2 A = pi*5^2 ... If the diameter is 10, then the radius is 5 A = 25pi So the area is 25pi square inches

OpenStudy (anonymous):

My Sister Says Thanks (:

jimthompson5910 (jim_thompson5910):

you're welcome